Subject: Minutes of APO 3.5m User's Committee Meeting, March 8, 1999

From: strauss@astro.Princeton.EDU

Submitted: Tue, 9 Mar 1999 12:00:05 -0500 (EST)

Message number: 340 (previous: 339, next: 341 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
        *******************March 8, 1999**********************

Attending: Michael Strauss, Rene Walterbos, Bruce Gillespie, Ed
Turner, Jeff Brown, Don York, Alan Uomoto, Ed Kibblewhite
Not attending: Chris Stubbs

******************The Status of the Echelle****************************
  York: There is little new news about the echelle since message 335
to apo35-general.  The commissioning time has been very productive.
Recent fixes include some software fixes and tracking down scattered
light.  The echelle is operable, and simple to operate.  One can guide
remotely; this is very simple to do.  Uses 1.6" slit, which gives
proper sampling on the CCD.  The main limitation is the read noise of
the chip: 7 e-.  One can do 1 hour exposures in practice; much more
than this gives unacceptable cosmic ray rates. With this, one is
photon-limited brighter than 14 mag.  Cosmic Ray rate seems a bit
higher than other chips on the mountain.  Throughput numbers haven't
changed; we don't know exactly why it is as low as it is.  (The
throughput is 3% from outside the atmosphere; the spectrograph itself
delivers 6% assuming nominal reflectance of the telescope optics,
compared to the 9% design value).  Gillespie, with input from Stubbs,
will carry out careful measurements, with SPICAM, of the throughput of
the telescope itself.

  The coverage is 3100-9500 A, without any gaps (all orders fit on the
  The guide camera does fine at 16th mag on a dark night (remember
that the full moon sky is 16 mag/square arcsec). 
  Software is in decent shape; IRAF standard reductions do a full job
of reducing these data. 
  There is some tilt of the orders with respect with CCD rows.  If one
were to do 2x2 binning to beat the read noise, you might get some
aliasing between orders. 
   We have observed a large range of objects: quasars, stars, etc. 
   We are still looking for subtle problems, but haven't found any
   The stability seems to be very good, and the data are straightforward to
calibrate (a single wavelength calibration per night should be
enough).  There are no moving parts in the instrument under normal
operations.  A number of people have observed with it, to tease out
problems in observing software, etc. 
  All calibration lamps are built in, and are properly light-tight. 
  The MC is not yet capable of running two instruments simultaneously; so
you can't yet calibrate the echelle, while someone else is observing
with another instrument (but see below).
  Documentation/user's manual will exist by April 1, when this becomes
a user instrument, as well as a maintenance document for the mountain
  Observing with echelle on-site is very straightforward.  For remote
observing, people will want to talk to Don York or Doug Duncan to get
a 1/2 hour tutelage over the phone.  The first person from each
institution to use the instrument can then teach others at the site. 
But we would advise people to go to the site on their first

The web site for the echelle can be found at:
Check it out to see some representative spectra. 

  Gillespie: service observing with the echelle:  Because the echelle
can observe bright objects in poor conditions, it can be used (for
service observing) when other instruments cannot be used.  This will
make lots of sense once the tertiary rotation is automated.  This may
be the first steps towards queue observing at APO. 

******************Status of the secondary****************************
  Uomoto: They have finished polishing the mirror.  They found some
astigmatism in the final tests.  They believe (hope?) that this is due
not to the secondary, but rather to the test setup.  They may figure
this out in the next few days.
******************The New Top End****************************
  Kibblewhite: I am starting on a strawman design with high-bandwidth
tip-tilt system.  The system will be light-weight; we need to think
about thermal control, mechanical stiffness, and cost.  I am talking
to Jon Davis, and the folks involved in the 2.5m system.  I will have
the first design ready for discussion at the next user's committee

******************New Infrared Imaging Camera*********************
  Turner: Gillespie and I have been talking to Sean Casey about
getting his 1KX1K InSb infrared imaging instrument on as a visiting
instrument.  Field of view is roughly 1 arcminute.  Lots of enthusiasm
all around.  Sean is in the process of writing up a "white paper", or
draft MOU, about how to carry this out.  The relevant time frame is
something like 6 months.   

  Gillespie: GRIM continues to show signs of age; it is becoming more
unreliable.  Sandford is working on the electronics, improving its

Gillespie: Loomis and Lowenstein have been working on the remark/MC
interface in the context of getting the echelle working.  This should
lead to the point that one should be able to operate multiple
instruments at a single time, on the timescale of a month or so.  

  The vBNS link through Las Cruces now hooks every one of the APO
institutions (except Washington State) to the site.  We've seen
substantial improvements in latency.  But there is nothing but
apocryphal reports about how improved things are.  Send your reports
of experiences to Gillespie.

  There has been a slight shift of the position of the primary,
investigated by Jon Davis.  There seems to be no stress point on the
mirror.  This introduces a small amount of astigmatism in the image.
People have been reporting slightly elliptical images in focus.  This
is a non-trivial thing to fix; Jon Davis is suggesting doing this
during the summer shut-down, when he would improve a number of things
about the primary support system.

There was amechanical accident last night: one of intermediate-level
hatch doors was left open in the afternoon.  Telescope enclosure was
moved, and ran into this door, destroying it, and screwing up the
enclosure mechanism.  Assessment is happening now.  Enclosure drive
wheels appear to be fine.  The telescope itself was not affected.  It
is not yet clear whether it will be possible to observe tonight. 

  Camron Hastings will be staying through July, giving us some
breathing room for finding a replacement for him. 

  Seeing monitor is in place; we'll wait for Chris Stubbs for a full

  Tertiary rotation will be commissioned during 19 April through 2 May
(tentative dates only!). 

  Last month's minutes are approved. 

  Next meeting: April 12, 12 noon. 

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 340 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to